OpenStudy (anonymous):

7x+5 over 4 = x+2 and 9(6+2a) = -(2-4a)

OpenStudy (gorv):

\[7x +\frac{ 5 }{ 4 }=x+2\]

OpenStudy (gorv):

subtract x from both side

OpenStudy (gorv):

7x-x+5/4=2 6x+5/4=2

OpenStudy (gorv):

subtract 5/4 from both side and after this divide both side by 6

OpenStudy (gorv):

9*6+9*2a=-2+4a 54=18a=-2+4a

OpenStudy (gorv):

54+18a=-2+4a

OpenStudy (anonymous):

7x + 5 = 4x + 2 3x = -3 x= -1 54 + 18a = -2 + 4a 14 a = -56 a = -4

OpenStudy (gorv):

subtract 4a from both side and subtract 54 from both side